# Tractable and Provably Efficient Distributional Reinforcement Learning with General Value Function Approximation

### v1

@@ 4 Statistical Functionals in Distributional RL
4.1 Bellman Unbiasedness
We revisit the definitions of statistical functionals and illustrate the property of Bellman closedness described in Bellemare et al. [9]. Then we introduce another pivotal property, Bellman unbiasedness, required for learning in terms of statistical functionals, rather than approximated distribution.
Definition 4.1 (Statistical functionals, Sketch; [9]).
@@ 4.1 Bellman Unbiasedness
We denote the domain of sketch as $\mathscr{P}_{\psi_{1:N}}(\mathbb{R})$ and its image as $I_{\psi_{1:N}}=\{\psi(\bar{\eta}):\bar{\eta}\in\mathscr{P}_{\psi_{1:N}}(\mathbb{R})\}.$ We further extend to state return distribution functions $\psi_{1:N}(\bar{\eta})=\Big{(}\psi_{1:N}(\bar{\eta}(s)):s\in\mathcal{S}\Big{)}.$
Definition 4.2 (Bellman closedness; [41]).
@@
which is closed under the distributional Bellman operator $\mathcal{T}:\mathscr{P}(\mathbb{R})^{\mathcal{S}}\rightarrow\mathscr{P}(\mathbb{R})^{\mathcal{S}}$.
Bellman closedness is the property that a sketch are exactly learnable when updates are performed from the infinite-dimensional distribution space to the finite-dimensional embedding space. Notably, Rowland et al. [41] showed that the only finite linear statistical functionals that are Bellman closed are given by the collections of statistical functionals where its linear span is equal to the set of exponential-polynomial functionals where $\psi_{0}$ is the constant functional equal to $1$.
Remark 4.3.
Rowland et al. [41] discussed only linear statistical functionals when defining a sketch, leaving open questions about the Bellman closedness of nonlinear statistical functionals such as variance, central moments, or quantiles. However, it is noteworthy that nonlinear statistical functionals, such as maximum or minimum, can also be Bellman closed. In their proof, there might be an ambiguity regarding the assumption that the corresponding sketch Bellman operator for quantile is linear. In this paper, we show the non-existence of sketch Bellman operator for quantile and discuss in detail in Appendix B.1.
@@
We investigate whether there exists another sketch that is both nonlinear and Bellman closed, and introduce additional conditions under which the Bellman closed property might be missing. During the implementation phase, it is important to note that the agent does not have access to the transition kernel. Specifically, the agent can only access the empirical transition kernel $\hat{\mathbb{P}}$ which is derived from a finite number of sample estimates. This limitation implies that the operator should be treated as an empirical operator $\hat{\mathcal{T}}_{\psi}$, rather than $\mathcal{T}_{\psi}$. Therefore, we introduce a new notion of Bellman unbiasedness that allows us to unbiasedly estimate the expected distribution $(\mathcal{B}_{r})_{\#}\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}[\bar{\eta}(s^{\prime})]$, which is a mixture by transitions, from the sample distribution $(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})$.
${\eta}$${\mathcal{T}\eta}$${\psi(\mathcal{T}\eta)}$${\psi(\eta)}$${\hat{\mathcal{T}}_{\psi}\psi(\eta)}$${{\mathbb{E}_{\mathbb{P}}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi_{\psi}\circ}\hat{\mathcal{T}}_{\psi}\psi(\eta)]}}$  Sample  next state Unbiasedly estimate
Figure 3: (Left) Bellman Closedness. (Right) Bellman Unbiasedness. The above path represents an ideal distributional Bellman update. Due to the infinite-dimensionality, the update process should be represented by using a finite-dimensional embedding (sketch) $\psi$. Since the transition kernel $\mathbb{P}$ is unknown, the below path describes that the implementation should sample the next state and update by using $\hat{\mathcal{T}}_{\psi}$ with the empirical transition kernel $\hat{\mathbb{P}}$. A sketch $\psi$ is Bellman unbiased if $\hat{\mathcal{T}}_{\psi}\circ\psi$ can unbiasedly estimate $\psi\circ\mathcal{T}$ through some transformation $\phi_{\psi}$.
@@
A sketch ${\psi}$ is Bellman unbiased if there exists an integer $k$ and a corresponding vector-valued estimator $\phi_{\psi}=\phi_{\psi}(\psi(x_{1}),\cdots,\psi(x_{k})):(I_{\psi}^{\mathcal{S}})^{k}\rightarrow I_{\psi}^{\mathcal{S}}$ where the sketch of expected distribution $(\mathcal{B}_{r})_{\#}\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}[\bar{\eta}(s^{\prime})]$ can be unbiasedly estimated by $\phi_{\psi}$ using the $k$ samples from the sketch of the sample distribution $(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})$, i.e.,
$\mathbb{E}_{s^{\prime}_{1},\cdots,s^{\prime}_{k}\sim\mathbb{P}(\cdot|s,a)}\Bigg{[}{\phi_{\psi}}\Bigg{(}\underbrace{\psi\Big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime}_{1})}_{\hat{\mathcal{T}}_{\psi}\psi(\eta)}\Big{)},\cdots,\psi\Big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s_{k}^{\prime})\Big{)}\Bigg{)}\Bigg{]}$ $=\psi\Big{(}(\mathcal{B}_{r})_{\#}\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}[\bar{\eta}(s^{\prime})]\Big{)}.$
Bellman unbiasedness is another natural definition, similar to Bellman closedness, which takes into account a finite number of samples for the transition. As a simple example, median functional is not Bellman unbiased as the sample median is not an unbiased estimate of the median functional. In general, there is no unbiased estimator for the median, implying that the median functional cannot be unbiasedly estimated within an embedding space represented by a finite number of statistical functionals. On the other hand, variance functional is a nonlinear statistical functional but can be unbiasedly estimated. Then, the following question naturally arises;
“Which sketches are unbiasedly estimatable under the sketch-based Bellman update?”
@@
Lemma 4.6 states that in statistical functional dynamic programming, the unbiasedly estimatable embedding of a distribution can only be structured in the form of functions that are homogeneous of finite degree. (e.g., variance functional is a homogeneous of degree $2$ [23]). Taking this concept further and combining it with the results on Bellman closedness, we prove that even when including a nonlinear statistical functional, the only sketch that can be unbiasedly estimated in a finite-dimensional embedding space is the moment functional.
@@
Compared to Rowland et al. [41], our results extend beyond linear statistical functionals to include unbiasedly estimatable nonlinear statistical functionals, showing the uniqueness of the moment functional. Our theoretical results not only show that high-order central moments such as variance or skewness are exactly learnable, but also reveal that other nonlinear statistical functionals like median or quantiles inevitably involve approximation errors due to biased estimations. In Section 4.2, we emphasize that unbiasedly estimable sketches play an important role in mitigating misspecification errors, thereby enabling us to achieve sublinear regret.
@@ 4.2 Statistical Functional Bellman Completeness
We consider distributional reinforcement learning with value function approximation. For successful TD learning, reinforcement learning with function approximation commonly requires the assumption, Bellman Completeness, that after applying Bellman operator, the output lies in the function class $\mathcal{F}$ [50, 4, 25]. On the other hand, our approach receives a tuple of function class $\mathcal{F}^{N}\subseteq\{f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}\}$ as input to represent $N$-moment of distribution. Building on this, we assume that for any $\bar{\eta}:\mathcal{S}\rightarrow\mathscr{P}([0,H])$, the sketch of target function that results from applying sketch Bellman operator lies in the function class $\mathcal{F}^{N}$.
In the seminal works, Wang et al. [49] and Chen et al. [11] assumed that the function class $\mathcal{H}\subseteq\{\eta:\mathcal{S}\times\mathcal{A}\rightarrow\mathscr{P}([0,H])\}$ follows the distributional Bellman Completeness assumption, i.e., if $\eta\in\mathcal{H}$ for all $\pi,h\in[H]$, then $\mathcal{T}_{h}^{\pi}\eta\in\mathcal{H}$. However, since the distributional Bellman operator mixes distributions for the next state transition, it implies that a function class $\mathcal{H}$ must be closed under mixture. Due to this closedness under mixture, both previous studies assumed a discretized reward MDP, which requires the prior knowledge that outcomes of return distribution are additively closed. In general, representing a distribution with a fixed number of parameters inevitably increases the required number of parameters due to the emergence of a mixture distribution. Consequently, trying to represent it with a limited number of parameters leads to approximating the learning outcome, thereby causing model misspecification which generally introduces an additional linear regret. Precisely, a regret lower bound with model misspecification error $\zeta$ is known as $\Omega(\zeta K)$ in a linear bandit setting [52]. Therefore, assuming Bellman Completeness on the distribution space is invalid as it significantly constrains the MDP structure or causes linear regret.
To circumvent this issue, we revisit the assumption of distributional Bellman completeness through the statistical functional lens. We propose a new framework that matches a finite number of all statistical functionals to the target function, rather than the entire distribution itself.
@@
Assumption 4.9 (Model Misspecification).
There exists a set of function $\mathcal{F}^{N}$ and a real number $\zeta>0$, such that for any $\bar{\eta}:\mathcal{S}\rightarrow\mathscr{P}([0,H])$ and $h\in[H]$, there exists $f_{\bar{\eta}}\in\mathcal{F}^{N}$ which satisfies
$\max_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left\|f_{\bar{\eta}}(s,a)-\psi_{1:N}\big{(}(\mathcal{B}_{r_{h}})_{\#}[\mathbb{P}_{h}\bar{\eta}](s,a)\big{)}\right\|\leq\zeta\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$
We call $\zeta$ the misspecification error.
5 Algorithms
In this section, we propose SF-LSVI for distributional RL framework with general value function approximation. Leveraging the result from Theorem 4.7, we introduce a moment least square regression. This allows us to capture a finite set of moment information from the distribution, which can be unbiasedly estimated, thereby leading to the truncated moment problem. Unlike previous work [49, 11] that estimates in infinite-dimensional distribution spaces, our method enables tractable distribution estimation in finite-dimensional embedding spaces without approximation error.
[…]
1:  for episode $k=1,2,\dots,K$ do
2:   Receive initial state $s_{1}^{k}$
3:   Initialize $\psi_{1:N}(\bar{\eta}_{H+1}^{k}(\cdot))\leftarrow\bm{0}^{N}$
4:   for step $h=H,H-1,\dots,1$ do
11:    $\psi_{1}\Big{(}\bar{\eta}^{k}_{h}(\cdot,\cdot)\Big{)}\leftarrow V^{k}_{h}(\cdot)$,  $\psi_{2:N}\Big{(}\bar{\eta}^{k}_{h}(\cdot)\Big{)}\leftarrow\psi_{1:N}\Big{(}\eta^{k}_{h}(\cdot,\pi^{k}_{h}(\cdot))\Big{)}_{n\in[2:N]}$
12:   for $h=1,2,\dots,H$ do
13:    Take action $a^{k}_{h}\leftarrow\pi^{k}_{h}(s^{k}_{h})$
14:    Observe reward $r^{k}_{h}(s^{k}_{h},a^{k}_{h})$ and get next state $s^{k}_{h+1}$.

### v2

@@ 3 Statistical Functionals in Distributional RL
In this section, we define two key concepts in the distRL framework: the statistical functional and the sketch. We also illustrate Bellman closedness, a crucial property from Bellemare et al. (2023). Next, we introduce Bellman unbiasedness, a novel concept that complements the previous property and is essential for provable efficiency. As shown in Figure 2, quantile functionals cannot be updated in an unbiased manner (as proved in Theorem 3.3), demonstrating that only certain sketches can be updated exactly. We then show that the only sketch satisfying both properties is the moment functional, which is unique among statistical functionals. Finally, we discuss the intractability of the previous structural assumption, distributional Bellman Completeness, and its tendency to cause linear regret. To address this, we introduce statistical functional Bellman Completeness, a relaxed assumption, and explain why it satisfies both properties.
Figure 3: Bellman Closedness
${\eta}$${\mathcal{T}\eta}$${\psi(\mathcal{T}\eta)}$${\psi(\eta)}$${\hat{\mathcal{T}}_{\psi}\psi(\eta)}$${{\mathbb{E}_{\mathbb{P}}[{\phi_{\psi}\circ}\hat{\mathcal{T}}_{\psi}\psi(\eta)]}}$  Sample  next state Unbiasedly estimate
Figure 4: Bellman Unbiasedness
Figure 5: Illustration of Bellman Closedness and Bellman Unbiasedness. The above path represents an ideal distributional Bellman update. Due to the infinite-dimensionality, the update process should be represented by using a finite-dimensional embedding (sketch) $\psi$. Since the transition kernel $\mathbb{P}$ is unknown, the below path describes that the implementation should sample the next state and update by using $\hat{\mathcal{T}}_{\psi}$ with the empirical transition kernel $\hat{\mathbb{P}}$. A sketch $\psi$ is Bellman unbiased if $\hat{\mathcal{T}}_{\psi}\circ\psi$ can unbiasedly estimate $\psi\circ\mathcal{T}$ through some transformation $\phi_{\psi}$, i.e., $\psi(\mathcal{T}\eta)=\mathbb{E}_{\mathbb{P}}[\phi_{\psi}\circ\hat{\mathcal{T}}\psi(\eta)]$.
3.1 Bellman Closedness
Definition 3.1 (Statistical functionals, Sketch; (Bellemare et al., 2023)).
@@ 3.1 Bellman Closedness
We denote the domain of sketch as $\mathscr{P}_{\psi_{1:N}}(\mathbb{R})$ and its image as $I_{\psi_{1:N}}=\{\psi_{1:N}(\bar{\eta}):\bar{\eta}\in\mathscr{P}_{\psi_{1:N}}(\mathbb{R})\}.$ We further extend to state return distribution functions $\psi_{1:N}(\bar{\eta})=\Big{(}\psi_{1:N}(\bar{\eta}(s)):s\in\mathcal{S}\Big{)}$
Definition 3.2 (Bellman closedness; (Rowland et al., 2019)).
@@
which is closed under a distributional Bellman operator $\mathcal{T}:\mathscr{P}(\mathbb{R})^{\mathcal{S}}\rightarrow\mathscr{P}(\mathbb{R})^{\mathcal{S}}$.
Bellman closedness is the property that a sketch are exactly learnable when updates are performed from the infinite-dimensional distribution space to the finite-dimensional embedding space. While classical Bellman equation implies the existence of Bellman operator for expected values, not all statistical functional has such corresponding Bellman operator. Precisely, Rowland et al. (2019) showed that the only finite linear statistical functionals that are Bellman closed are given by the collections of statistical functionals where its linear span is equal to the set of exponential-polynomial functionals.  [^4] In discounted setting, a unique solution becomes moments. We’ve overwritten it for convinience.
@@
While Rowland et al. (2019) focused on "linear" statistical functionals in defining a sketch (i.e., $\psi(\bar{\eta})=\mathbb{E}_{Z\sim\bar{\eta}}[h(Z)]$ for some $h$), leaving questions about nonlinear functionals, we extend this by showing that "nonlinear" statistical functionals, such as maximum or minimum, can also be Bellman closed. Additionally, while their proof implicitly treated quantiles as linear functionals, we provide a technical clarification in Appendix C.1 where we formally demonstrate that no sketch Bellman operator exists for quantiles.
3.2 Bellman Unbiasedness
While the intractability caused by infinite-dimensionality was addressed in Bellman closedness, another intractable element which has not yet fully tackled is the sampling of the next state. During the implementation, note that the agent does not have access to the transition kernel $\mathbb{P}$. Instead, the agent can only access the empirical transition kernel ^$\hat{\mathbb{P}}(\cdot|s,a)=\frac{1}{K}\sum_{k=1}^{K}\bm{1}\{s^{\prime}_{k}=\cdot\ |s,a\}$ which is derived from $K$ sampled next states. This limitation implies that the operator should be treated as an empirical operator $\hat{\mathcal{T}}_{\psi}$, rather than $\mathcal{T}_{\psi}$ (i.e., $\hat{\mathcal{T}}_{\psi}\psi(\bar{\eta})\coloneqq\psi((\mathcal{B}_{r})_{\#}[\hat{\mathbb{P}}\bar{\eta}])$). Therefore, we naturally introduce a new notion of Bellman unbiasedness to unbiasedly estimate the expected distribution $(\mathcal{B}_{r})_{\#}\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}[\bar{\eta}(s^{\prime})]$, which is a mixture by transitions, from the sample distribution $(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})$.
@@ 3.2 Bellman Unbiasedness
A sketch ${\psi}(=\psi_{1:N})$ is Bellman unbiased if a vector-valued estimator $\phi_{\psi}=\phi_{\psi}(\psi(\cdot),\cdots,\psi(\cdot)):(I_{\psi}^{\mathcal{S}})^{k}\rightarrow I_{\psi}^{\mathcal{S}}$ exists where the sketch of expected distribution $(\mathcal{B}_{r})_{\#}\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}[\bar{\eta}(s^{\prime})]$ can be unbiasedly estimated by $\phi_{\psi}$ using the $k$ sampled sketches from the sample distribution $(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime})$, i.e.,
$\mathbb{E}_{s^{\prime}_{i}\sim\mathbb{P}}\Bigg{[}{\phi_{\psi}}\Bigg{(}\underbrace{\psi\Big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s^{\prime}_{1})\Big{)},\cdots,\psi\Big{(}(\mathcal{B}_{r})_{\#}\bar{\eta}(s_{k}^{\prime})\Big{)}}_{{k\text{ sampled sketches from sample distribution }}\hat{\mathcal{T}}_{\psi}\psi(\bar{\eta}(s))}\Bigg{)}\Bigg{]}$
$=\psi\Big{(}(\mathcal{B}_{r})_{\#}\mathbb{E}_{s^{\prime}\sim\mathbb{P}(\cdot|s,a)}[\bar{\eta}(s^{\prime})]\Big{)}.$
Bellman unbiasedness is another natural definition, similar to Bellman closedness, which takes into account a finite number of samples for the transition. For example, mean-variance sketch is Bellman unbiased as the following unbiased estimator $\phi_{(\mu,\sigma^{2})}$ exists for $k$ sample estimates:
$(\mu,\sigma^{2})$ $=\phi_{(\mu,\sigma^{2})}\Big{(}(\hat{\mu}_{1},\hat{\sigma}_{1}^{2}),\cdots,(\hat{\mu}_{k},\hat{\sigma}_{k}^{2})\Big{)}$
$=\Big{(}\frac{1}{k}\sum_{i=1}^{k}\hat{\mu}_{i},\ \frac{1}{k}\sum_{i=1}^{k}(\hat{\mu}_{i}-\frac{1}{k}\sum_{i=1}^{k}\hat{\mu}_{i})^{2}+\hat{\sigma}_{i}^{2}\Big{)}$
On the other hand, median functional is not Bellman unbiased since there is no unbiased estimator for median. Then, the following question naturally arises;
Which sketches are unbiasedly estimatable under the sketch-based Bellman update?
@@
Lemma 3.5 states that in statistical functional dynamic programming, the unbiasedly estimatable embedding of a distribution can only be structured in the form of functions that are homogeneous of finite degree (Halmos, 1946). To illustrate that homogenity defines a broader class than linear functionals, consider the variance as a simple example. Variance is clearly not a linear linear functional, as it is non-additive. However, it can be written as
$\text{Var}(\bar{\eta})=\mathbb{E}_{Z_{1}\sim\bar{\eta}}[(Z_{1}-\mathbb{E}_{Z_{2}\sim\bar{\eta}}[Z_{2}])^{2}]$
$=\mathbb{E}_{Z_{1},Z_{2}\sim\bar{\eta}}[Z_{1}^{2}-2Z_{1}Z_{2}+Z_{2}^{2}]=\mathbb{E}_{Z_{1},Z_{2}\sim\bar{\eta}}[h(Z_{1},Z_{2})]$
which implies the homogenity of degree $2$. Taking this concept further and combining it with the results on Bellman closedness, we prove that even when including a nonlinear statistical functional, the only sketch that can be exactly learned and unbiasedly estimated in a finite-dimensional embedding space is the moment sketch.
@@
Compared to Rowland et al. (2019), we extend beyond linear statistical functionals to include nonlinear statistical functionals, showing the uniqueness of the moment functional. As shown in Figure 1, our theoretical results not only show that high-order central moments such as variance or skewness are exactly learnable and unbiasedly estimatable, but also reveal that other nonlinear statistical functionals like median or quantiles inevitably involve approximation errors due to biased estimations.
Necessity of Bellman unbiasedness.
Bellman unbiasedness ensures that updates can be unbiasedly performed when only a finite number of sampled sketches are available. In other words, it guarantees that the sequence of sampled sketches forms a martingale, enabling the construction of confidence regions through concentration inequalities. This property is crucial for establishing provable efficiency in terms of regret minimization.
Complementary roles of unbiasedness and closedness.
At first glance, Bellman Unbiasedness (BU) may appear to be a stricter subset of Bellman Closedness (BC). However, as illustrated in Figure 1, the relationship is more subtle: for example, the categorical sketch is BU but not BC, whereas functionals like the maximum or minimum are BC but not BU. More precisely, BU guarantees the existence of an unbiased estimator of the ground-truth sketch given a finite number of sampled sketches. In contrast, BC plays a complementary role by ensuring that the update process consistently provide such sketches. If a sketch is BU but not BC–as in the case of the categorical sketch–then the update process cannot continue providing new sampled sketches, making dynamic programming infeasible.
@@ 3.3 Statistical Functional Bellman Completeness
We consider distributional reinforcement learning with general value function approximation (GVFA). For successful TD learning, GVFA framework for classical RL commonly requires the assumption, Bellman Completeness, that after applying Bellman operator, the output lies in the function class $\mathcal{F}$ (Wang et al., 2020; Ayoub et al., 2020; Ishfaq et al., 2021). As a natural extension, our approach receives a tuple of function class $\mathcal{F}^{N}\subseteq\{f:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{N}\}$ as input to represent $N$ moments of distribution. Building on this, we assume that for any $\bar{\eta}:\mathcal{S}\rightarrow\mathscr{P}([0,H])$, the sketch of target function lies in the function class $\mathcal{F}^{N}$.
@@
DistBC inevitably leads to linear regret.
In the seminal works, Wang et al. (2023) and Chen et al. (2024) assumed that the function class $\mathcal{H}\subseteq\{\eta:\mathcal{S}\times\mathcal{A}\rightarrow\mathscr{P}([0,H])\}$ follows the distributional Bellman Completeness (distBC) assumption (i.e., if $\eta\in\mathcal{H}$ for all $\pi,h\in[H]$, $\mathcal{T}_{h}^{\pi}\eta\in\mathcal{H}$). This seems natural, but constructing a finite-dimensional subspace $\mathcal{H}$ that satisfies distBC is quite challenging. Since the distributional Bellman operator is a composition of translation and mixing distributions for the next state, it implies that a function class $\mathcal{H}$ must be closed under translation and mixture. However, when considering the representation of infinite-dimensional distributions using a finite number of representations, it is not trivial that the mixture of distributions can also be represented with the same number of representations. For example, while a Gaussian distribution can be represented using two parameters $(\mu,\sigma^{2})$, a mixture of $K$ Gaussians generally requires $2K$ representations.
To avoid the issue of closedness under mixture, both previous studies assumed a discretized reward MDP where all outcomes of the return distribution are able to discretized into an uniform grid of finite points. Unfortunately, the approximation error introduced by the discretization is not negligible when it comes to regret. This is because model misspecification, which is the error when the model fails to represent the target, typically leads to linear regret.
Definition 3.8 (Model Misspecification in distBC).
For a given distribution class $\mathcal{H}$ which is the finite-dimensional subspace of the space of all distribution $\mathcal{F}^{\infty}$, we call $\zeta$ the misspecification error
$\zeta\coloneqq\inf_{f_{\bar{\eta}}\in\mathcal{H}}\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\|f_{\bar{\eta}}(s,a)-(\mathcal{B}_{r_{h}})_{\#}[\mathbb{P}_{h}\bar{\eta}](s,a)\|$
for any $\bar{\eta}:\mathcal{S}\rightarrow\mathscr{P}([0,H])$ and $h\in[H]$.
Note that $\zeta$ is strictly positive unless the function approximator $f_{\bar{\eta}}$ can represent any distribution in the finite-dimensional subspace $\mathcal{H}$ generated by translation and mixture. In a classical linear bandit setting (Zanette et al., 2020), a lower bound with misspecification error $\zeta$ is known to yield linear regret $\Omega(\zeta K)$. Therefore, redefining Bellman Completeness within the infinite-dimensional distribution space is not appropriate, as it either imposes strong constraints on the MDP structure or leads to linear regret. To circumvent model misspecification, we revisit the distributional BC through the statistical functional lens. We propose a novel framework that matches a finite number of statistical functionals to the target, rather than the entire distribution itself.
4 SF-LSVI: Statistical Functional Least Square Value Iteration
In this section, we propose SF-LSVI for distRL framework with general value function approximation. Leveraging the result from Theorem 3.6, we introduce a moment least square regression. This allows us to capture a finite set of moment information from the distribution, which can be unbiasedly estimated, thereby leading to the truncated moment problem (Shohat & Tamarkin, 1943; Schmüdgen et al., 2017). Unlike previous work (Wang et al., 2023; Chen et al., 2024) that estimates in infinite-dimensional distribution spaces, our method enables to estimate distribution unbiasedly in finite-dimensional embedding spaces without misspecification error. The pseudocode of SF-LSVI is described in Appendix B.